# Unleashing the Potential of Tracklets for Unsupervised Video Person Re-Identification

### v1

@@ Abstract
With rich temporal-spatial information, video-based person re-identification methods have shown broad prospects. Although tracklets can be easily obtained with ready-made tracking models, annotating identities is still expensive and impractical. Therefore, some video-based methods propose using only a few identity annotations or camera labels to facilitate feature learning. They also simply average the frame features of each tracklet, overlooking unexpected variations and inherent identity consistency within tracklets. In this paper, we propose the Self-Supervised Refined Clustering (SSR-C) framework without relying on any annotation or auxiliary information to promote unsupervised video person re-identification. Specifically, we first propose the Noise-Filtered Tracklet Partition (NFTP) module to reduce the feature bias of tracklets caused by noisy tracking results, and sequentially partition the noise-filtered tracklets into “sub-tracklets”. Then, we cluster and further merge sub-tracklets using the self-supervised signal from tracklet partition, which is enhanced through a progressive strategy to generate reliable pseudo labels, facilitating intra-class cross-tracklet aggregation. Moreover, we propose the Class Smoothing Classification (CSC) loss to efficiently promote model learning. Extensive experiments on the MARS and DukeMTMC-VideoReID datasets demonstrate that our proposed SSR-C for unsupervised video person re-identification achieves state-of-the-art results and is comparable to advanced supervised methods.
Index Terms:
Person re-identification, Unsupervised learning, Video tracklets, Reliable clustering
@@ I Introduction
Person Re-IDentification (Re-ID) aims to retrieve a query identity from a gallery of disjoint cameras [1, 2]. In recent years, person Re-ID has made great progress with the prosperity of deep neural networks. Compared with image-based methods [3, 4, 5, 6, 7, 8, 9, 10, 11, 12], video-based methods [13, 14, 15, 16, 17, 18, 19, 20] usually achieve better performance with the help of rich temporal-spatial information. Nonetheless, prevailing video-based approaches predominantly rely on supervised learning, necessitating costly cross-camera tracklet annotations, thereby impeding their broader deployment in practical real-world settings. Given the substantial challenges and impracticalities associated with obtaining a wealth of annotated tracklets, alternative strategies have emerged: (1) One-shot based methods [21, 22, 23, 24] propose to label only one tracklet for each identity; (2) Tracklet association based methods [25, 26, 17] explore the correspondence of within-camera and cross-camera tracklets, which assumes the camera labels are available. However, these methods still require several pedestrian tracklet annotations or camera information, thereby restricting their universal applicability in diverse real-world scenarios.
@@
Drawing inspiration from established practices in unsupervised image-based Re-ID methods [27, 28, 29, 30, 31], a more practical and feasible solution is to perform the clustering algorithm to generate pseudo identity labels for training. One of the most direct ways is to conduct clustering directly based on the tracklet features. However, since tracklets are obtained using off-the-shelf tracking models in an unsupervised manner, noisy frames caused by tracking errors are inevitable, and some difficult frames, e.g. with heavy occlusions or appearance variations, can also introduce bias to the tracklet features. On the one hand, as illustrated in Fig. 1 (a), with the tracklet overlapping between pedestrians, the tracklet features of different individuals may be close in the embedding space, so that different identities are incorrectly assigned the same pseudo label. On the other hand, as shown in Fig. 1 (b), due to the presence of occlusions and background variations, tracklets of the same identity may be mistakenly assigned to different clusters. The noises and unexpected variations of tracklet features would weaken the representativeness and discriminability of tracklets, undermining their capacity to accurately depict distinct identities and reducing their effectiveness in distinguishing between identities. As a result, imprecise clustering results can lead to noisy pseudo labels to misguide the learning of the model. Additionally, we notice that the noises within a tracklet and the similarity between tracklets exhibit locality. For example, the overlapping of the two pedestrians only exists in some of their whole tracklets in Fig. 1 (a), and there are similar visual features in local tracklets in Fig. 1 (b). It is worth exploring and taking advantage of this locality to help refine clustering results, promoting effective feature learning.
@@
Therefore, we first propose the Noise-Filtered Tracklet Partition (NFTP) module. It filters noisy frames within each tracklet in an adaptive manner and then sequentially partition each tracklet into multiple segments, referred to as “sub-tracklets”. Based on noise-filtered sub-tracklets, we propose the self-supervised refined clustering to generate pseudo labels effectively. To improve the quality of generated pseudo labels, it first acquires reliable sub-clusters via performing clustering under strict restrictions, and then merge sub-clusters into ultimate clusters. More specifically, we propose to make use of the identity consistency between sub-tracklets, that is, sub-tracklets from the same original tracklet share the same identity, to merge sub-clusters. In this process, due to the inherent local temporal-spatial similarity between tracklets, noise-filtered sub-tracklets would act as the intra-class relays to aggregate different cross-camera tracklets of the same pedestrian. As our method does not rely on annotations or auxiliary information, it inevitably encounters clustering inaccuracies, which can exert a notable impact on the model’s learning process. Of particular concern is that during the early stages of training, the model’s poor feature extraction capabilities may exacerbate the prevalence of such clustering mistakes. Additionally, as training goes on, the model becomes more robust and the extracted features become more reliable. Therefore, we propose a progressive merging strategy to alleviate the impact of inevitable clustering inaccuracies on model learning, which improves the discriminative ability and robustness of the model effectively. Eventually, with reliable pseudo labels and sub-tracklets with less feature bias, we propose the Class Smoothing Classification (CSC) loss to effectively guide the model learning at a more fine-grained sub-tracklet level. By iteratively performing our proposed self-supervised refined clustering and effective feature learning, the capability and robustness of the model can be significantly enhanced. Our contributions can be summarized as follows:
@@ II-A Unsupervised Video Person Re-Identification
Tracklet association based methods tend to investigate tracklet association across different cameras [26, 25, 17, 20]. For example, Li et al. [25] simultaneously conduct tracklet association learning within-camera and across-camera, as well as a soft labeling approach to mitigate penalties for potentially positive samples. Chen et al. [26] explore local space-time consistency within each tracklet from the same camera view and global cyclic ranking consistency between tracklets across disjoint camera views. Zang et al. [17] also explore the potential of local-level features for unsupervised learning. However, the reliance on additional camera information renders them inflexible in real-world applications where such information may be inaccessible.
Clustering-based methods provide a more effective way for unsupervised person Re-ID [32, 33, 34, 35, 36, 37, 38, 39]. They adopt clustering for pseudo identity labels, which are used for model optimization. Fan et al. [32] propose a progressive unsupervised learning method that iterates between pedestrian clustering with K-means [40] and fine-tuning the model. However, it requires a pre-defined number of clusters, which can be impractical in real-world applications. Bottom-Up Clustering (BUC) [33] progressively incorporates samples into clusters, but negative samples can be inevitably included, impacting the quality of the generated pseudo labels. Following BUC, some works alleviate the adverse impact of noisy frames during clustering [35] or introduce a dispersion-based criterion to achieve high-quality clustering [36]. Wu et al. [34] design a comprehensive unsupervised learning objective that accounts for tracklet frame coherence, tracklet neighborhood compactness, and tracklet cluster structure. Xie et al. [38] resample and re-weight the hard frames in tracklets to improve clustering ability. However, due to the inherent suboptimal clustering robustness of these methods, it is difficult for them to achieve satisfactory performance.
@@ II-B Self-Supervised Learning for Person Re-Identification
Figure 2: The framework of our proposed SSR-C. During each training epoch, our proposed Noise-Filtered Tracklet Partition (NFTP) module would filter noisy frames within each tracklet, and further sequentially partition each tracklet into multiple sub-tracklets. Different shapes (e.g., triangle and circle) correspond to different tracklets, different colors (e.g., red and blue) denote different identities, and various transparencies of the same color represent the sub-tracklets of each tracklet. With noise-filtered sub-tracklets, we can acquire reliable sub-clusters via strictly restricted clustering, and then sub-clusters are merged leveraging the self-supervised signal from tracklet partition and following a progressive merging strategy. Finally, we design the Class Smoothing Classification (CSC) loss, which leverages the obtained accurate pseudo labels to effectively improve the discriminative ability of the model.
III Proposed Method
III-A Clustering-based Framework Overview
@@
We first provide the overview of the clustering-based baseline that is based on the original tracklets. During the training process, given an unlabeled training set $X=\left\{T_{1},T_{2},\cdots,T_{N}\right\}$ of $N$ tracklets, where each tracklet contains $L_{i}$ frames, i.e., $T_{i}=\left\{x_{i,j}\right\}_{j=1}^{L_{i}}$. $M$ frames of each tracklet are randomly selected and average pooled to figure out the feature embedding $v_{i}$ for each tracklet. Formally,
$v_{i}=\frac{1}{M}\sum_{j=1}^{M}\phi\left(\theta;x_{i,j}\right)$ (1)
@@
where $v_{i}\in\mathbb{R}^{d}$, and $\phi$ is the feature embedding model parameterized by $\theta$. For notation simplicity, we denote $\phi\left(\theta;x_{i,j}\right)$ as $f_{i,j}\in\mathbb{R}^{d}$. Since the identity annotations of tracklets are inaccessible under the unsupervised setting, a density-based clustering method DBSCAN [45] is performed to obtain pseudo labels, which does not require prior knowledge of the total number of identities.
@@
After applying DBSCAN, tracklets clustered together are assigned to the same pseudo label and isolated outliers are simply discarded. Then, we can get a “labeled” dataset $\hat{X}=\{\left(T_{i},\hat{y_{i}}\right)\}_{i=1}^{N}$, where $1\leq\hat{y_{i}}\leq n$ denotes the pseudo label of the $i$-$th$ tracklet, and $n$ is the total number of clusters. Additionally, we calculate the centroid of each cluster by averaging tracklet features within it to initialize the memory $\boldsymbol{V}\in\mathbb{R}^{n\times d}$. Based on the pseudo labels and the memory, InfoNCE [46] loss is adopted for the model training. Formally,
$\mathcal{L}_{c}=-\log p(\hat{y}_{i}\mid T_{i},\boldsymbol{V})=-\log\frac{\exp\left(\boldsymbol{v_{i}}^{\top}\cdot\boldsymbol{V}_{\hat{y_{i}}}/\tau\right)}{\sum_{j=1}^{n}\exp\left(\boldsymbol{v_{i}}^{\top}\cdot\boldsymbol{V}_{j}/\tau\right)}$ (2)
@@
where $\boldsymbol{V}_{j}$ is the $j$-th entry of $\boldsymbol{V}$, and $\tau$ is the temperature parameter [47] to control the softness of probability distribution over classes. During back-propagation, we update $\boldsymbol{V}$ with the momentum update strategy as follows:
$\boldsymbol{V}_{j}\leftarrow\alpha\boldsymbol{V}_{j}+\left(1-\alpha\right)\overline{v}_{j}$ (3)
@@
where $\overline{v}_{j}$ is the average of tracklet features with the pseudo label $j$ in this mini-batch.
@@
To further distinguish easily confused sample pairs and improve the inter-class separability, similar to [48], we also maintain an additional hard memory $\boldsymbol{V}^{h}$ storing the hard tracklet feature of each class. $\boldsymbol{V}^{h}$ is updated in a similar way as Eq. 3, except that we mine the hardest sample in the mini-batch for the hard memory update. Specifically, we compute the cosine similarity between each sample with pseudo label $\hat{y}_{i}$ in the mini-batch and the corresponding memory entry $\boldsymbol{V}_{\hat{y_{i}}}^{h}$, and the hardest sample for class $\hat{y}_{i}$ is the one with the lowest cosine similarity. The overall loss function is formulated as:
$\mathcal{L}_{h}=-\log\left(p\left(\hat{y}_{i}\mid T_{i},\boldsymbol{V}^{h}\right)\right)$ (4)
$\mathcal{L}=\gamma_{1}\mathcal{L}_{h}+\gamma_{2}\mathcal{L}_{c}$ (5)
@@
where $\gamma_{1}$ and $\gamma_{2}$ are hyper-parameters to balance between $\mathcal{L}_{h}$ and $\mathcal{L}_{c}$. By iteratively performing clustering and model training, the feature extraction capability of the model as well as the accuracy of the obtained pseudo labels could be gradually enhanced.
III-B Noise-Filtered Tracklet Partition

### v2

@@ Abstract
With rich temporal-spatial information, video-based person re-identification methods have shown broad prospects. Although tracklets can be easily obtained with ready-made tracking models, annotating identities is still expensive and impractical. Therefore, some video-based methods propose using only a few identity annotations or camera labels to facilitate feature learning. They also simply average the frame features of each tracklet, overlooking unexpected variations and inherent identity consistency within tracklets. In this paper, we propose the Self-Supervised Refined Clustering (SSR-C) framework without relying on any annotation or auxiliary information to promote unsupervised video person re-identification. Specifically, we first propose the Noise-Filtered Tracklet Partition (NFTP) module to reduce the feature bias of tracklets caused by noisy tracking results, and sequentially partition the noise-filtered tracklets into “sub-tracklets”. Then, we cluster and further merge sub-tracklets using the self-supervised signal from the tracklet partition, which is enhanced through a progressive strategy to generate reliable pseudo labels, facilitating intra-class cross-tracklet aggregation. Moreover, we propose the Class Smoothing Classification (CSC) loss to efficiently promote model learning. Extensive experiments on the MARS and DukeMTMC-VideoReID datasets demonstrate that our proposed SSR-C for unsupervised video person re-identification achieves state-of-the-art results and is comparable to advanced supervised methods. The code is available at https://github.com/Darylmeng/SSRC-Reid.
@@ I Introduction
Person Re-IDentification (Re-ID) aims to retrieve a query identity from a gallery of disjoint cameras [1, 2]. In recent years, person Re-ID has made great progress with the prosperity of deep neural networks [3, 4, 5, 6, 7, 8, 9]. Compared with image-based methods [10, 11, 12, 13, 14], video-based methods [15, 16, 17, 18, 19, 20, 21, 22] usually achieve better performance with the help of rich temporal-spatial information. Nonetheless, prevailing video-based approaches predominantly rely on supervised learning, necessitating costly cross-camera tracklet annotations, thereby impeding their broader deployment in practical real-world settings. Given the substantial challenges and impracticalities associated with obtaining a wealth of annotated tracklets, alternative strategies have emerged: (1) One-shot based methods [23, 24, 25, 26] propose to label only one tracklet for each identity; (2) Tracklet association based methods [27, 28, 19, 22] explore the correspondence of within-camera and cross-camera tracklets, which assumes the camera labels are available. However, these methods still require several pedestrian tracklet annotations or camera information, thereby restricting their universal applicability in diverse real-world scenarios.
@@
Drawing inspiration from established practices in unsupervised image-based Re-ID methods [29, 30, 31, 32, 33], a more practical and feasible solution is to perform the clustering algorithm to generate pseudo identity labels for training. The straightforward implementation is to conduct clustering directly based on the tracklet features. However, since tracklets are obtained using off-the-shelf tracking models in an unsupervised manner, noisy frames caused by tracking errors are inevitable, and some difficult frames, e.g., with heavy occlusions or appearance variations, can also introduce bias to the tracklet features.
On the one hand, as illustrated in Fig. 1 (a), with the tracklet overlapping between pedestrians, the tracklet features of different individuals may be close in the embedding space, so that different identities are incorrectly assigned the same pseudo label. On the other hand, as shown in Fig. 1 (b), due to the presence of occlusions and background variations, tracklets of the same identity may be mistakenly assigned to different clusters. The noises and unexpected variations of tracklet features would weaken the representativeness and discriminability of tracklets, undermining their capacity to accurately depict distinct identities and reducing their effectiveness in distinguishing between identities. As a result, imprecise clustering results can lead to noisy pseudo labels to misguide the learning of the model. Additionally, we notice that the noises within a tracklet and the similarity between tracklets exhibit locality. For example, the overlapping of the two pedestrians only exists in some of their whole tracklets in Fig. 1 (a), and there are similar visual features in local tracklets in Fig. 1 (b). It is worth exploring and taking advantage of this locality to help refine clustering results, promoting effective feature learning.
@@
Therefore, we first propose the Noise-Filtered Tracklet Partition (NFTP) module. It filters noisy frames within each tracklet in an adaptive manner and then sequentially partition each tracklet into multiple segments, referred to as “sub-tracklets”. Based on the noise-filtered sub-tracklets, we propose the self-supervised refined clustering to generate pseudo labels effectively. To improve the quality of generated pseudo labels, we first acquire reliable sub-clusters via performing clustering under strict restrictions, and then merge sub-clusters into ultimate clusters. More specifically, we make use of the identity consistency between sub-tracklets, that is, sub-tracklets from the same original tracklet share the same identity, to merge sub-clusters. In this process, due to the inherent local temporal-spatial similarity between tracklets, noise-filtered sub-tracklets would act as the intra-class relays to aggregate different cross-camera tracklets of the same pedestrian.
As our method does not rely on annotations or auxiliary information, it inevitably encounters clustering inaccuracies, which can exert a notable impact on the model’s learning process. Of particular concern is that during the early stages of training, the model’s poor feature extraction capabilities may exacerbate the prevalence of such clustering mistakes. Additionally, as training goes on, the model becomes more robust and the extracted features become more reliable. Therefore, we propose a progressive merging strategy to alleviate the impact of inevitable clustering inaccuracies on model learning, which improves the discriminative ability and robustness of the model effectively. Eventually, with reliable pseudo labels and sub-tracklets with less feature bias, we propose the Class Smoothing Classification (CSC) loss to effectively guide the model learning at a more fine-grained sub-tracklet level. By iteratively performing our proposed self-supervised refined clustering and effective feature learning, the capability and robustness of the model can be significantly enhanced. Our contributions can be summarized as follows:
@@ II-A Unsupervised Video Person Re-Identification
II-A1 Tracklet association based methods
Of this type, research tends to investigate tracklet association across different cameras [28, 27, 19, 22]. For example, Li et al. [27] simultaneously conduct tracklet association learning within-camera and across-camera, as well as a soft labeling approach to mitigate penalties for potentially positive samples. Chen et al. [28] explore local space-time consistency within each tracklet from the same camera view and global cyclic ranking consistency between tracklets across disjoint camera views. Zang et al. [19] also explore the potential of local-level features for unsupervised learning. However, the reliance on additional camera information renders them inflexible in real-world applications where these clues may be inaccessible.
II-A2 Clustering-based methods
These methods provide a more effective way for unsupervised person Re-ID [34, 35, 36, 37, 38, 39, 40, 41, 42]. They adopt clustering for pseudo identity labels, which are used for model optimization. Fan et al. [34] propose a progressive unsupervised learning method that iterates between pedestrian clustering with K-means [43] and fine-tuning the model. However, it requires a pre-defined number of clusters, which can be impractical in real-world applications. Bottom-Up Clustering (BUC) [35] progressively incorporates samples into clusters, but negative samples can be inevitably included, impacting the quality of the generated pseudo labels. Following BUC, some works alleviate the adverse impact of noisy frames during clustering [37] or introduce a dispersion-based criterion to achieve high-quality clustering [38]. Wu et al. [36] design a comprehensive unsupervised learning objective that accounts for tracklet frame coherence, tracklet neighborhood compactness, and tracklet cluster structure. Dai et al. [40] employ a memory dictionary and computes contrastive loss at the cluster level. Xie et al. [41] resample and re-weight the hard frames in tracklets to improve clustering ability. Tao et al. [44] perform a denoising diffusion process [45] and guide global feature learning in patch context to refine pseudo labels. However, due to the inherent suboptimal clustering robustness of these methods, it is difficult for them to achieve satisfactory performance.
@@ II-B Self-Supervised Learning for Person Re-Identification
III Methodology
We first formulate the unsupervised video person Re-ID task in Sec. III-A. Then, we provide the overview of the clustering-based baseline that is based on the original tracklets in Sec. III-B, where we extend the image-based work [50] to the video scenario as our baseline. Considering harmful factors within tracklets, we propose noise-filtered tracklet partition strategy, which results in sub-tracklets (See Sec. III-C). Based on sub-tracklets, we further introduce self-supervised refined clustering in Sec. III-D, resulting in abundant pseudo labels of sub-tracklets. After that, to optimize the model effectively, we introduce an effective loss in Sec. III-E. Finally, we introduce the training details of our proposed clustering-based framework in Sec. III-F.
III-A Task Formulation
Given an unlabeled training set $X=\left\{T_{1},T_{2},\cdots,T_{N}\right\}$ of $N$ tracklets, where each tracklet contains $L_{i}$ frames, i.e., $T_{i}=\left\{x_{i,j}\right\}_{j=1}^{L_{i}}$, the goal is to optimize the model $\phi$ to extract discriminative identity feature from each input tracklet $T_{i}$. During inference, the optimized model $\phi$ is used to extract features of tracklets from the query and gallery sets, respectively. Then, the features of the query tracklets are used to search the gallery set and retrieve the matched tracklets according to the cosine distance metric.
@@ III-A Task Formulation
Figure 2: The framework of our proposed SSR-C. During each training epoch, our proposed Noise-Filtered Tracklet Partition (NFTP) module would filter noisy frames within each tracklet, and further sequentially partition each tracklet into multiple sub-tracklets. Different shapes (e.g., triangle and circle) correspond to different tracklets, different colors (e.g., red and blue) denote different identities, and various transparencies of the same color represent the sub-tracklets of each tracklet. With noise-filtered sub-tracklets, we can acquire reliable sub-clusters via strictly restricted clustering, and then sub-clusters are merged leveraging the self-supervised signal from tracklet partitioning and following a progressive merging strategy. Finally, we design the Class Smoothing Classification (CSC) loss, which leverages the obtained accurate pseudo labels to effectively improve the discriminative ability of the model.
III-B Clustering-based Framework Overview
III-B1 Pseudo Label Acquisition
@@
During the training process, $M$ frames of each tracklet are randomly selected and average pooled to figure out the feature embedding $\boldsymbol{v}_{i}$ for each tracklet. Formally,
$\boldsymbol{v}_{i}=\frac{1}{M}\sum_{j=1}^{M}\phi\left(\theta;x_{i,j}\right),$ (1)
@@
where $\boldsymbol{v}_{i}\in\mathbb{R}^{d}$, and $\phi$ is the feature embedding model parameterized by $\theta$. For notation simplicity, we denote $\phi\left(\theta;x_{i,j}\right)$ as $f_{i,j}\in\mathbb{R}^{d}$. Since the identification annotations of the tracklets are inaccessible in the unsupervised setting, a density-based clustering method DBSCAN [51] is performed to obtain pseudo labels, which does not require prior knowledge of the total number of identities.
@@
After applying DBSCAN, the clustered tracklets are assigned to the same pseudo label and isolated outliers are simply discarded. Then, we can get a “labeled” dataset $\hat{X}=\{\left(T_{i},\hat{y_{i}}\right)\}_{i=1}^{N}$, where $1\leq\hat{y_{i}}\leq n$ denotes the pseudo label of the $i$-th tracklet, and $n$ is the total number of clusters. Additionally, we calculate the centroid of each cluster by averaging tracklet features within it to initialize a memory $\boldsymbol{V}\in\mathbb{R}^{n\times d}$.
III-B2 Training
At each training step, based on the pseudo labels and the memory, InfoNCE [52] loss is adopted for model optimization. Formally,
$\mathcal{L}_{c}$ $=-\log\left(p\left(\hat{y}_{i}\mid T_{i},\boldsymbol{V}\right)\right)$ (2)
$=-\log\frac{\exp\left(\boldsymbol{v_{i}}^{\top}\cdot\boldsymbol{V}_{\hat{y_{i}}}/\tau\right)}{\sum_{j=1}^{n}\exp\left(\boldsymbol{v_{i}}^{\top}\cdot\boldsymbol{V}_{j}/\tau\right)},$ (3)
@@ III-B2 Training
where $\boldsymbol{v}_{i}$ and $\hat{y_{i}}$ are the feature embedding and pseudo label of the $i$-th tracklet, $\boldsymbol{V}_{j}$ is the $j$-th entry of $\boldsymbol{V}$, and $\tau$ is the temperature parameter [53] to control the softness of the probability distribution over classes. During back-propagation, we update $\boldsymbol{V}$ with the momentum update strategy as follows:
$\boldsymbol{V}_{j}\leftarrow\alpha\boldsymbol{V}_{j}+\left(1-\alpha\right)\overline{\boldsymbol{v}}_{j},$ (4)
@@
where $\overline{v}_{j}$ is the average of tracklet features with the pseudo label $j$ in the mini-batch.
@@
To further distinguish easily confused sample pairs and improve the inter-class separability, similar to [50], we also maintain an additional hard memory $\boldsymbol{V}^{h}$ storing the hard tracklet feature of each class. $\boldsymbol{V}^{h}$ is initialized the same as $\boldsymbol{V}$ and is updated in a similar way to Eq. 4, except that we mine the hardest sample in the mini-batch for the hard memory update. Specifically, we compute the cosine similarity between each sample with pseudo label $\hat{y}_{i}$ in the mini-batch and the corresponding memory entry $\boldsymbol{V}_{\hat{y_{i}}}^{h}$, and the hardest sample for class $\hat{y}_{i}$ is the one with the lowest cosine similarity. The overall loss function is formulated as:
$\mathcal{L}_{h}=-\log\left(p\left(\hat{y}_{i}\mid T_{i},\boldsymbol{V}^{h}\right)\right),$ (5)
$\mathcal{L}=\gamma_{1}\mathcal{L}_{h}+\gamma_{2}\mathcal{L}_{c},$ (6)
@@
where Eq. 5 is defined similar to Eq. 3, $\gamma_{1}$ and $\gamma_{2}$ are hyper-parameters to balance between $\mathcal{L}_{h}$ and $\mathcal{L}_{c}$. By iteratively performing clustering and model training, the feature extraction capability of the model as well as the accuracy of the obtained pseudo labels could be gradually enhanced.
III-C Noise-Filtered Tracklet Partition